OpenStudy (anonymous):

Show that a line through (0, 0) and (c, d) is perpendicular to a line through (0, 0) and (-d, c). Slope of first line = Slope of second line =

OpenStudy (anonymous):

well if a slope is y/x then a perpendicular one is -x/y

OpenStudy (anonymous):

soit would b 0?

OpenStudy (anonymous):

so d/c firls line -c/d second line

OpenStudy (anonymous):

wrong answers.. those are not it

OpenStudy (anonymous):

tried those already

OpenStudy (anonymous):

you say sow that a line through (0, 0) and (c, d) is perpendicular to a line through (0, 0) and (-d, c). i am showing that :P

OpenStudy (anonymous):

:/ i already tried sthose... :c

OpenStudy (anonymous):

well you arer right :)

OpenStudy (anonymous):

Show that a line through (0, 0) and (c, d) is perpendicular to a line through (0, 0) and (-d, c). Slope of first line = Slope of second line = --------------- well line through (0, 0) and (c, d) has slope (d-0)/(c-0)=d/c a line through (0, 0) and (-d, c). has slope (c-0)/(-d-0)=-(c/d) therfore d/c was inverted and multiplied by -1 to make it perpendicular

OpenStudy (anonymous):

all you need to do to show that two lines are perpendicular is multiply their slopes... if it equals -1, you're done... if it does not, then they are not perpendicular.

OpenStudy (anonymous):

wait so what do i put :/

OpenStudy (anonymous):

the rules is to just take a slope invert it and multiply by negative one y/x -x/y 1/2 -2/1 3/5 -5/3 d/c - c/d because well line through (0, 0) and (c, d) has slope (d-0)/(c-0)=d/c a line through (0, 0) and (-d, c). has slope (c-0)/(-d-0)=-(c/d)